OpenStudy (anonymous):

182 divided by 6 = ?

OpenStudy (anonymous):

30.33333.....

OpenStudy (anonymous):

what else do you want done here?

OpenStudy (anonymous):

You should try to simplify it: 182/6 = 91/3 = 30 Remainder 1. Or 30.3